### v2

# Automatic Classification of General Movements in Newborns

Abstract
General movements (GMs) are spontaneous, coordinated body movements in infants that offer valuable insights into the developing nervous system. Assessed through the Prechtl GM Assessment (GMA), GMs are reliable predictors for neurodevelopmental disorders. However, GMA requires specifically trained clinicians, who are limited in number. To scale up newborn screening, there is a need for an algorithm that can automatically classify GMs from infant video recordings. This data poses challenges, including variability in recording length, device type, and setting, with each video coarsely annotated for overall movement quality. In this work, we introduce a tool for extracting features from these recordings and explore various machine learning techniques for automated GM classification.
keywords:
Automated General Movements Assessment (GMA), Neurodevelopmental Disorder Screening Tool, Time Series, Paediatrics
Data and Code Availability
The private dataset consists of infant videos collected by the Barmherzige Brüder Regensburg Hospital, details in Section 2.1 and Appendix A. The code [^1] https://github.com/chopardda/GMA is publicly available.
Institutional Review Board (IRB)
IRB project 23-3304-104: “Development of a Computer tool for the Assessment of General Movements”.
1 Introduction and Related Work
General movements (GMs) refer to spontaneous movements of the entire body observable in infants from early fetal life until approximately six months post-term (Einspieler and Prechtl, 2005). These movements include coordinates sequences of the arms, legs, neck, and trunk, with variations in intensity, force, and speed. GMs are inherently complex, variable, and fluid, and their quality can provide critical insights into the integrity of the developing nervous system. Assessing GM quality, known as the Prechtl General Mouvement Assessment (GMA), is a sensitive and reliable diagnostic tool for predicting the likelihood of cerebral palsy and other neurodevelopmental disorders (Einspieler and Prechtl, 2005). The types of movements predictive of later neurodevelopmental disorders vary significantly with age, and age-based distinctions are essential when studying newborns. From early fetal life until approximately two months post-term, GMs exhibit consistent writhing movements. Between 6 to 9 weeks post-term, these writhing patterns diminish, and fidgety movements emerge, persisting until around the middle of the first year of life, at which point intentional movements become predominant. Alterations in these movement patterns can indicate later disorders. While GMA is typically conducted in hospitals by specialized clinicians, there are too few trained experts to screen all newborns. This motivates the need to develop a tool for automatic and reliable classification of movements.
Initial attempts at addressing this challenge focused on computer-based approaches using feature extraction techniques (Baccinelli et al., 2020; Raghuram et al., 2022). However, recent advances in deep learning and computer vision have spurred the development of more sophisticated methods (Silva et al., 2021; Irshad et al., 2020). For instance, Reich et al. (2021) proposed to use a pose estimator, OpenPose, to extract skeletons and used these as input to a shallow multilayer neural network to discriminate between movements. Other deep learning solutions, e.g. by Schmidt et al. (2019), work in controlled conditions. Despite their promise, these approaches often rely on controlled environments with fixed sensors and constant recording length, limiting their scalability and applicability in real-world clinical settings.
In this work, we present a preliminary study introducing a method to label key anatomical points, automatically process videos, and classify GMs based on the tracked anatomical points across videos, using a dataset from Barmherzige Brüder Regensburg Hospital. The challenges posed by the variability in recording lengths and devices, the diversity of video scenarios (including both hospital and home environments), and the nature of the movement quality annotations—one per recording—are significant for reliable GM classification. Addressing these challenges is crucial to the generalization of an automatic GMA that can be effectively deployed in varied settings. The goal is to enable comprehensive newborn screening and follow-up studies by experts, ultimately improving the early detection and treatment of neurodevelopmental disorders.
2 Dataset and Methods
This work involves labelling specific anatomical landmarks of interest—so-called keypoints (KP)—in video frames (Section 2.2), tracking their positions across the video (Section 2.2), and extracting features from the tracked keypoints (Section 2.4) to classify movement quality. The full pipeline is illustrated in Figure 1.
Figure 1: Flowchart of keypoint extraction pipeline after initial video resampling and cropping.
2.1 Dataset
This study includes 76 infant video recordings collected from the Barmherzige Brüder Regensburg Hospital. The dataset is highly heterogeneous, including videos recorded both at home and in hospital settings using various recording devices, resulting in differing resolutions and frame rates. There is also significant variability in camera angles and orientations, distance from subject to camera, infants’ clothing, and background types, among other factors.
The data is split into two age-based groups: the early infancy group includes 39 infants with a postmenstrual age of up to 37 weeks, while the late infancy group consists of 37 infants with postmenstrual ages ranging from 46 to 56 weeks. Each sample is associated with a binary label that classifies the infant’s movement quality. In the early group, labels distinguish between normal and poor movement repertoire, while in the late group, labels indicate normal or absence of fidgety movements, following Einspieler and Prechtl (2005). To minimize bias and improve reliability, two trained physicians independently annotated the videos. In cases of disagreement, the recordings were reviewed jointly until a consensus label was reached. In the dataset, 24 infants in the early group exhibit poor repertoire, while 10 in the late group lack fidgety movements.
2.2 Keypoint Retrieval Tools
Keypoint Labelling
We introduce a keypoint labelling tool based on the work from Doersch et al. (2022) to mark sets of key coordinates in the videos, which, along with point tracking, forms the backbone of the preprocessing pipeline. The tool allows labelling of extremities such as wrists and ankles—referred to as extreme keypoints, as well as additional joints such as elbows, hips, and knees—collectively referred to as all keypoints, totaling 17. In this study, three independent non-expert labellers manually labelled all keypoints for subsequent tracking. Full lists of relevant keypoints can be found in Appendix B, see Figure 2 a) for an illustration of the tool.
Point Tracking
For tracking labelled keypoints across video frames, we utilize the point tracking algorithm proposed by Doersch et al. (2022). This approach predicts the trajectory of each point by extracting dense feature representations for every frame, leveraging a cost-volume approach. The cost volume measures the similarity between the feature of the tracked point in one frame and the features at all possible locations in subsequent frames, encoding the probability of the tracked point’s next position. A neural network then processes this cost volume to predict the tracked point’s location across frames. More details are provided in Appendix C.
Figure 2: a) The point labelling tool prompts the user to label either only extreme keypoints or all keypoints. b) The outlier detection tool displays potential unrealistic point movements and allows manual correction.
2.3 Data Preprocessing
Due to the heterogeneity of the data, a robust preprocessing and error correction pipeline is necessary to prepare the videos for machine learning model development. This subsection outlines the key steps in that process, with a graphical overview in Figure 1.
Resampling and Extreme Keypoints Labelling
Preprocessing begins with homogenizing all the video data by resampling to 30 frames per second. Extreme keypoints are then labelled. If all required keypoints are not visible in the first frame, labelling occurs in subsequent frames where visibility is clearer.
Video Cropping
Next, the extracted extreme keypoints are tracked and used to apply a rectangular crop per video. The crop is determined by the most extreme values of the tracked points in each video, with a 15% margin to ensure the infant consistently occupies a similar portion of the video frame. This normalization step is crucial for maintaining consistency in the input data for machine learning models.
Outlier Detection
To address inaccuracies in the tracking algorithm, such as when points jump inappropriately due to occlusions or overlapping limbs, an outlier detection mechanism is applied. The tool automatically flags any sudden and unrealistic movement of keypoints and prompts manual relabelling by the annotators, as shown in Figure 2 b). After relabelling, the affected keypoints are retracked. This process is iterated a fixed number of rounds, two in these results, or until no further outliers are detected, ensuring the accuracy of the tracked coordinates.
Final Data Preparation
Prior to classification, the time series data must be homogenized. All video samples are split into fixed-length fragments corresponding to the shortest present recording, and zero-filling is applied to any occluded or missing coordinates. To prevent data leakage, we apply stratified sampling, ensuring that no single subject appears in both the training and testing datasets.
2.4 Feature Extraction
We consider two feature sets from the tracked keypoints: $x$- and $y$-coordinates of all keypoints and the angles between selected keypoints, based on Prakash et al. (2023) and further described in Appendix E.
Keypoint Coordinates
The $x$- and $y$-coordinates of all labelled keypoints are treated independently as separate input channels, and used as a simple representation of the keypoint positions across frames.
Keypoint Angles
The angles are calculated between pairs of keypoints to capture movement dynamics. We compute the dot product of vectors, normalizing by magnitudes, and applying a clipping function for valid angle calculations in radians.
Table 1: AUROC of various classifiers for GM classification across different keypoint labellers. The mean and standard deviation over 30 seeds is reported. Best result per model and labeller is bolded.
Keypoint Labeller Input Features Early Infancy Group AUROC Late Infancy Group AUROC
RF LSTM CNN RF LSTM CNN
Coord. $0.6574\pm 0.19$ $0.6398\pm 0.16$ $0.7630\pm 0.19$ $0.3678\pm 0.25$ $0.4449\pm 0.23$ $\textbf{0.5804}\pm\textbf{0.23}$
#1 Angles $\textbf{0.8283}\pm\textbf{0.13}$ $0.6617\pm 0.16$ $\textbf{0.8185}\pm\textbf{0.14}$ $\textbf{0.5125}\pm\textbf{0.25}$ $\textbf{0.6068}\pm\textbf{0.21}$ $0.3770\pm 0.22$
Both $0.7878\pm 0.14$ $\textbf{0.6655}\pm\textbf{0.16}$ $0.7876\pm 0.17$ $0.3763\pm 0.24$ $0.4007\pm 0.19$ $0.5371\pm 0.22$
Coord. $0.4907\pm 0.19$ $\textbf{0.6807}\pm\textbf{0.17}$ $0.654\pm 0.19$ $0.4538\pm 0.24$ $0.4906\pm 0.24$ $0.5096\pm 0.20$
#2 Angles $\textbf{0.6763}\pm\textbf{0.15}$ $0.6324\pm 0.17$ $0.6050\pm 0.19$ $\textbf{0.6284}\pm\textbf{0.20}$ $0.4902\pm 0.22$ $\textbf{0.6908}\pm\textbf{0.22}$
Both $0.5737\pm 0.20$ $0.6088\pm 0.16$ $\textbf{0.6729}\pm\textbf{0.20}$ $0.4899\pm 0.23$ $\textbf{0.5471}\pm\textbf{0.18}$ $0.5888\pm 0.20$
Coord. $0.5874\pm 0.21$ $0.6176\pm 0.18$ $0.7493\pm 0.20$ $0.5179\pm 0.27$ $0.4286\pm 0.27$ $0.6196\pm 0.19$
#3 Angles $\textbf{0.8105}\pm\textbf{0.12}$ $0.5621\pm 0.16$ $\textbf{0.8138}\pm\textbf{0.14}$ $\textbf{0.6267}\pm\textbf{0.22}$ $\textbf{0.5351}\pm\textbf{0.23}$ $\textbf{0.5915}\pm\textbf{0.24}$
Both $0.7278\pm 0.18$ $\textbf{0.662}\pm\textbf{0.21}$ $0.7846\pm 0.17$ $0.5651\pm 0.18$ $0.4762\pm 0.25$ $0.5392\pm 0.29$
Coord. $0.5785$ 0.6460 $0.7221$ $0.4465$ $0.4547$ 0.5699
Mean Angles 0.7717 $0.6187$ $0.7458$ 0.5892 0.5440 $0.5531$
Both $0.6964$ $0.6454$ 0.7484 $0.4771$ $0.4747$ $0.5550$
2.5 Classification Models
This work explores three different classification models for GMA, using the extracted time series described in Section 2.4. 5-fold stratified cross-validation (CV) is performed across models to ensure robustness.
1D Convolutional Neural Network (1D-CNN)
1D-CNNs are a variant of the more common 2D convolutional neural network, which apply convolution operations to 1D signals, making them well-suited for processing time-series data. These networks have been shown to perform well when working with smaller labelled datasets and for specific applications (Kiranyaz et al., 2021).
Long Short-Term Memory (LSTM)
LSTM networks (Hochreiter and Schmidhuber, 1997) are a type of recurrent neural network capable of exploiting the temporal aspect of our data by learning relevant context information from previously seen time points. LSTMs are ideal for handling time series and assessing whether longer temporal memory helps in recognizing and classifying infant’s general movements.
Random Forest (RF)
RFs (Breiman, 2001) are powerful discriminative classifiers that perform well on a variety of datasets with minimal tuning (Biau and Scornet, 2016). It is an ensemble method that builds multiple decision trees trained on a random subset of data and combines their outputs to improve prediction accuracy and reduce overfitting. Although they do not explicitly account for the temporal structure of the data, their efficiency and accuracy make them a strong candidate for our classification tasks.
3 Experiments and Results
3.1 Implementation details
We performed a grid search over the hyperparameter space for each of the three classifier types in Section 2.5, choosing the values that performed best in the evaluation. The final models were trained using these hyperparameters with 5-fold stratified CV, and the results from the held-out test set, with 20% of the data, were collected. The classification models used included a 1D-CNN with binary cross entropy (BCE) loss, learning rate of 0.00001, batch size 6, a fully-connected bottleneck with 150 features, and trained for 150 epochs. The LSTM used the BCE loss, learning rate 0.001, batch size 6, 3 layers, a hidden size of 64 and trained for 200 epochs. Finally, the random forest had 170 estimators. All experiments were run using 30 independent seeds on each individual set. Our models were implemented using pytorch v2.3.1 and scikit-learn v1.3.0. The final classification was evaluated using accuracy, Area Under the Receiver Operating Characteristic Curve (AUROC), and Area Under the Precision–Recall curve (AUPRC).
3.2 Results
Table 1 shows the results of the different classification models, for each labeller individually as well as averaged, measured in AUROC. See Appendix F for a comprehensive overview of AUPRC and accuracy in each task. They demonstrate a variety of outcomes across age groups and feature sets, with no clear pattern indicating that one method is superior. However, the use of angle features, either isolated or combined, proves useful. Additionally, the early group shows significantly better overall results, showing a big difference in the difficulty of the task across age groups. One of the key takeaways from this study is that despite the relatively small and heterogeneous dataset, our approach has shown success in automatically classifying GM in newborns, with average AUROCs of up to 0.8283 in certain scenarios, opening promising lines of research for GMA “in the wild”.
4 Conclusion and Future Outlook
In conclusion, our study demonstrates that we can successfully predict GMA with a promising AUROC of up to 0.8283, with performance generally better for the early group. This highlights the potential of using automated methods to assess infant motor behaviour. Despite the challenges posed by diverse data sources and recording conditions in the data, our results show that automated GMA is feasible in real-world clinical settings. This is a significant step forward in the early detection of neurodevelopmental disorders. This study lays the groundwork for further advancements, emphasizing the need for larger datasets and more refined models and feature extraction approaches to capture the data complexities. Ultimately, the goal is to enhance early prediction of neurodevelopmental diseases, offering a valuable tool for clinicians and ultimately improving patient care.
One of the key weaknesses of this study is the relatively small dataset, which limits the robustness and generalizability of our results, as indicated by relatively large standard deviations across folds. Moreover, the heterogeneity of the data makes the problem more challenging, while broadening its impact in the clinics. Expanding the dataset would allow us to explore more sophisticated methods, such as Multiple Instance Learning (MIL), which could introduce interpretability by making predictions on a per-episode basis. Additionally, using Vision Transformers to directly process the videos could better capture spatial and temporal patterns, further improving classification accuracy. Future work within the current dataset could focus on developing further features from the extracted keypoints and exploring other model classes to leverage the temporal structure in the task.
[…]
DC is funded from the Strategic Focal Area “Personalized Health and Related Technologies (PHRT)” grant #2021-911 of the ETH Domain (Swiss Federal Institutes of Technology), and SL and KC from the Swiss State Secretariat for Education, Research and Innovation (SERI), contract #MB22.00047. The authors would like to thank Ričards Marcinkevičs and Heike Leutheuser for their insights and discussions on the development of this manuscript.
Appendix A Extended Dataset Description
In this section we present additional details about our dataset to the best of our ability. Some details are not available to us, such as weight at video for early group infants. Table 2 shows several demographic characteristics, as well as details about the videos themselves for both the early and late groups. Figure 3 provides histograms for the distributions of both video resolution and FPS for the early and late video groups.
Table 2: Calculated demographic and video characteristics
Characteristic Early Group Late Group
Age at video recording (Days) (mean, stddev) 30.9, 19.9 170.4, 24.8
Post-menstrual age at video recording (Days) (mean, stddev) 231.5, 6.5 N/A
Corrected Age at video recording (Days) (mean, stddev) N/A 92.5, 17.4
Weight at video recording (g) (mean, stddev) N/A 4148.2, 830.7
Video Length (s) (mean, stddev) 409, 144 224, 93
FPS (min, max) 30, 120 30, 120
Extremely Preterm 16 15
Very Preterm 20 19
Moderate to Late Preterm 3 3
Not Preterm 0 0
Figure 3: Histograms showing the distributions of a) early infancy group video resolution, b) early infancy group video FPS, c) late infancy group video resolution, and d) late infancy group video FPS
Appendix B Keypoint Labelling Tool
Figure 5 illustrates the keypoints used in this study and the labelling tool developed.
Figure 4:
List of all and extreme keypoints considered in this study, coloured by label on the right. All Keypoints Extreme nose head bottom head top x left ear right ear left shoulder right shoulder left elbow x right elbow x left wrist x right wrist x left hip right hip left knee x right knee x left ankle x right ankle x
Figure 5: Example of frame with labelled keypoints.
Appendix C Keypoint Tracking
We adopt the work by Doersch et al. (2022) for keypoint tracking. Their model, TAP-Net, is designed to track arbitrary physical points on surfaces over video sequences. It is trained on a mix of synthetic and real-world datasets. The primary training datasets are synthetic, including TAP-Vid-Kubric, which contains simulated videos with physically realistic interactions between deformable and rigid objects, and TAP-Vid-RGB-Stacking, a robotic environment with teleoperated geometric shapes, focusing on challenges like tracking textureless and symmetric objects. These datasets provide perfect ground-truth point tracks, enabling TAP-Net to learn precise point tracking in complex, occlusion-prone scenarios. It uses a cost-volume-based architecture, inspired by optical flow methods, to track point motion across frames by comparing dense feature grids at query points. The network is trained with two objectives: (1) Point location regression, using Huber loss to estimate the query point’s position and ensure precise tracking, and (2) Occlusion classification, using cross-entropy loss to predict when the point is occluded, enabling the model to handle occlusion scenarios effectively. TAP-Net is validated and tested on real-world datasets like TAP-Vid-Kinetics, which includes diverse human actions and complex scenes, and TAP-Vid-DAVIS, which focuses on challenging segment and object tracking tasks.
Appendix D Outlier Detection
To mitigate potential errors from the tracking tool, we implemented a tool that automatically flags any sudden, unrealistic movements of keypoints. Specifically, we define such movements as coordinate changes that exceed 15 times the overall standard deviation for each keypoint. This threshold is set deliberately high to avoid flagging normal, natural movements, as well as fidgety movements that can inherently exhibit high magnitudes without indicating error.
Once flagged, the affected keypoints are relabeled and subsequently retracked. This process is iteratively applied up to three times, as we observed only marginal improvements in performance after three rounds in the final detection task.
In general, this outlier-detection process can be structured to run for a set number of iterations, until a defined performance tolerance is reached, or until no further outliers are identified, ensuring the accuracy of the tracked coordinates.
Appendix E Feature Extraction
In this work, we perform the classification task on a set of features extracted from the obtained keypoints. In particular, we study the classification on three subsets, as reported in our results:
  • Coord. The keypoint coordinates, that correspond to the x- and y-locations of all keypoints, as listed in Figure 5.
  • Angles The angles between keypoints of interest. The values of the angle corresponding to each keypoint pair are provided separately to the model.
    $\mathbf{p}_{1}$ $\mathbf{p}_{2}$ $\mathbf{p}_{3}$ Explanation
    head top nose head bottom Head top to neck angle
    right ear nose left ear Right to left ear angle
    left elbow left shoulder head bottom Head to left shoulder angle
    right elbow right shoulder head bottom Head to right shoulder angle
    left wrist left elbow left shoulder Left elbow angle
    right wrist right elbow right shoulder Right elbow angle
    left knee left hip left shoulder Left hip angle
    right knee right hip right shoulder Right hip angle
    left hip left knee left ankle Left knee angle
    right hip right knee right ankle Right knee angle
    Table 3: List of angle components used in the study with intuitive explanations.
    Figure 6: Illustration of the angle features. The dash lines show the angles that we compute in the set of angle features.
    Given keypoints $\mathbf{p}_{1}$, $\mathbf{p}_{2}$, and $\mathbf{p}_{3}$ with coordinates $(x_{1},y_{1})$, $(x_{2},y_{2})$ and $(x_{3},y_{3})$ respectively. We would like to compute the angle between $\mathbf{p}_{1}$ and $\mathbf{p}_{3}$ at middle point $\mathbf{p}_{2}$. Thus, we define
    $\mathbf{v}_{1}=\begin{bmatrix}x_{1}-x_{2}\\ y_{1}-y_{2}\end{bmatrix},\quad\mathbf{v}_{2}=\begin{bmatrix}x_{3}-x_{2}\\ y_{3}-y_{2}\end{bmatrix}.$
    Then, the angle $\theta$ between the two vectors can be computed as:
    $\theta=\arccos\left(\frac{\mathbf{v}_{1}\cdot\mathbf{v}_{2}}{\|\mathbf{v}_{1}\|\|\mathbf{v}_{2}\|}\right),$
    where $\mathbf{v}_{1}\cdot\mathbf{v}_{2}$ is the dot product of the vectors and $\|\mathbf{v}_{1}\|$ and $\|\mathbf{v}_{2}\|$ are the magnitudes (or norms) of the vectors.
    Table 3 provides the list of the studied angle sets, which keypoints they are formed by, and a brief intuitive explanation. The angles are illustrated in 6.
  • Both The combination of both coordinates and angles, shown as ”both” in the results.
Appendix F Further results
Table 4 and Table 5 show the results of the presented classification models, for each labeller individually as well as averaged, grouped by age and feature extraction method used, measured in AUPRC and accuracy, respectively. The best performing method per task is marked in bold and overall both metrics resemble the behaviour of the AUROCs reported in the main text in Table 1.
Table 4: Results. AUPRC over 30 seeds is reported. The best result for each model and labeller is bolded.
Labeller Features Early Infancy Group Late Infancy Group
AUPRC AUPRC
RF LSTM CNN RF LSTM CNN
Coord. $0.7733\pm 0.14$ $\textbf{0.7503}\pm\textbf{0.13}$ $0.8585\pm 0.11$ $0.2589\pm 0.14$ $0.3603\pm 0.22$ $\textbf{0.4627}\pm\textbf{0.23}$
1 Angles $\textbf{0.8856}\pm\textbf{0.09}$ $0.7346\pm 0.16$ $\textbf{0.8772}\pm\textbf{0.1}$ $\textbf{0.389}\pm\textbf{0.25}$ $\textbf{0.47}\pm\textbf{0.24}$ $0.266\pm 0.17$
Both $0.8575\pm 0.09$ $0.7359\pm 0.15$ $0.8547\pm 0.13$ $0.2879\pm 0.19$ $0.2938\pm 0.15$ $0.3983\pm 0.22$
Coord. $0.665\pm 0.14$ $\textbf{0.7798}\pm\textbf{0.14}$ $\textbf{0.799}\pm\textbf{0.12}$ $0.3142\pm 0.18$ $0.38\pm 0.27$ $0.4778\pm 0.24$
2 Angles $\textbf{0.7703}\pm\textbf{0.13}$ $0.7097\pm 0.17$ $0.6867\pm 0.16$ $\textbf{0.4674}\pm\textbf{0.25}$ $\textbf{0.3814}\pm\textbf{0.22}$ $\textbf{0.4959}\pm\textbf{0.27}$
Both $0.7088\pm 0.15$ $0.7105\pm 0.15$ $0.7955\pm 0.14$ $0.3523\pm 0.18$ $0.3804\pm 0.22$ $0.4638\pm 0.25$
Coord. $0.7259\pm 0.15$ $0.6986\pm 0.17$ $0.8522\pm 0.13$ $0.3593\pm 0.19$ $0.3405\pm 0.23$ $\textbf{0.5538}\pm\textbf{0.21}$
3 Angles $\textbf{0.8685}\pm\textbf{0.09}$ $0.6655\pm 0.16$ $0.85\pm 0.14$ $\textbf{0.4929}\pm\textbf{0.28}$ $\textbf{0.4194}\pm\textbf{0.24}$ $0.4061\pm 0.25$
Both $0.8121\pm 0.13$ $\textbf{0.7461}\pm\textbf{0.17}$ $\textbf{0.8579}\pm\textbf{0.12}$ $0.3603\pm 0.18$ $0.3838\pm 0.22$ $0.4269\pm 0.27$
Coord. $0.7214$ 0.7429 0.8366 $0.3108$ $0.3603$ 0.4981
Mean Angles 0.8415 $0.7033$ $0.8046$ 0.4498 0.4236 $0.3893$
Both $0.7928$ $0.7308$ $0.8360$ $0.3335$ $0.3527$ $0.4297$
Table 5: Results. Accuracy over 30 seeds is reported. The best result for each model and labeller is bolded.
Labeller Features Early Infancy Group Late Infancy Group
Accuracy Accuracy
RF LSTM CNN RF LSTM CNN
Coord. $0.5962\pm 0.16$ $0.59\pm 0.15$ $0.6946\pm 0.15$ $0.5759\pm 0.14$ $0.5793\pm 0.11$ $0.6133\pm 0.16$
1 Angles $\textbf{0.7266}\pm\textbf{0.13}$ $0.5978\pm 0.14$ $\textbf{0.7047}\pm\textbf{0.16}$ $\textbf{0.6305}\pm\textbf{0.17}$ $\textbf{0.6785}\pm\textbf{0.13}$ $0.5701\pm 0.13$
Both $0.6596\pm 0.15$ $\textbf{0.6444}\pm\textbf{0.16}$ $0.6849\pm 0.12$ $0.6073\pm 0.16$ $0.5809\pm 0.16$ $\textbf{0.6168}\pm\textbf{0.15}$
Coord. $0.4542\pm 0.17$ $\textbf{0.5973}\pm\textbf{0.15}$ $0.6219\pm 0.13$ $0.6385\pm 0.13$ $0.5856\pm 0.15$ $0.5603\pm 0.21$
2 Angles $\textbf{0.6518}\pm\textbf{0.14}$ $0.5594\pm 0.14$ $0.5568\pm 0.17$ $\textbf{0.6765}\pm\textbf{0.08}$ $0.6297\pm 0.14$ $\textbf{0.6432}\pm\textbf{0.13}$
Both $0.5268\pm 0.18$ $0.5871\pm 0.16$ $\textbf{0.6275}\pm\textbf{0.15}$ $0.6444\pm 0.13$ $\textbf{0.6418}\pm\textbf{0.14}$ $0.6113\pm 0.16$
Coord. $0.5548\pm 0.18$ $0.6125\pm 0.14$ $0.6764\pm 0.14$ $0.6652\pm 0.14$ $0.5955\pm 0.15$ $0.614\pm 0.19$
3 Angles $\textbf{0.6995}\pm\textbf{0.11}$ $0.5443\pm 0.15$ $\textbf{0.7508}\pm\textbf{0.12}$ $\textbf{0.6995}\pm\textbf{0.08}$ $0.6233\pm 0.16$ $0.638\pm 0.12$
Both $0.6186\pm 0.17$ $\textbf{0.6502}\pm\textbf{0.17}$ $0.7101\pm 0.14$ $0.6605\pm 0.09$ $\textbf{0.6505}\pm\textbf{0.17}$ $\textbf{0.6393}\pm\textbf{0.13}$
Coord. $0.5351$ $0.5999$ $0.6643$ $0.6265$ $0.5868$ $0.5959$
Mean Angles 0.6926 $0.5672$ $0.6708$ 0.6688 0.6438 $0.6171$
Both $0.6017$ 0.6272 0.6742 $0.6374$ $0.6244$ 0.6225

### v3

@@
-Automatic Classification of General Movements in Newborns
+Towards Scalable Newborn Screening: Automated General Movement Assessment in Uncontrolled Settings
 Abstract
@@
 In this work, we introduce a tool for extracting features from these recordings and explore various machine learning techniques for automated GM classification.
-keywords:
-Automated General Movements Assessment (GMA), Neurodevelopmental Disorder Screening Tool, Time Series, Paediatrics
-Data and Code Availability
-The private dataset consists of infant videos collected by the Barmherzige Brüder Regensburg Hospital, details in Section 2.1 and Appendix A.
-The code [^1] https://github.com/chopardda/GMA is publicly available.
-Institutional Review Board (IRB)
-IRB project 23-3304-104: “Development of a Computer tool for the Assessment of General Movements”.
-1 Introduction and Related Work
-General movements (GMs) refer to spontaneous movements of the entire body observable in infants from early fetal life until approximately six months post-term (Einspieler and Prechtl, 2005).
+1 Introduction
+General movements (GMs) refer to spontaneous movements of the entire body observable in infants from early fetal life until approximately six months post-term (Einspieler & Prechtl, 2005).
 These movements include coordinates sequences of the arms, legs, neck, and trunk, with variations in intensity, force, and speed.
 GMs are inherently complex, variable, and fluid, and their quality can provide critical insights into the integrity of the developing nervous system.
-Assessing GM quality, known as the Prechtl General Mouvement Assessment (GMA), is a sensitive and reliable diagnostic tool for predicting the likelihood of cerebral palsy and other neurodevelopmental disorders (Einspieler and Prechtl, 2005).
+Assessing GM quality, known as the Prechtl General Mouvement Assessment (GMA), is a sensitive and reliable diagnostic tool for predicting the likelihood of cerebral palsy and other neurodevelopmental disorders (Einspieler & Prechtl, 2005).
 The types of movements predictive of later neurodevelopmental disorders vary significantly with age, and age-based distinctions are essential when studying newborns.
@@
 Despite their promise, these approaches often rely on controlled environments with fixed sensors and constant recording length, limiting their scalability and applicability in real-world clinical settings.
-In this work, we present a preliminary study introducing a method to label key anatomical points, automatically process videos, and classify GMs based on the tracked anatomical points across videos, using a dataset from Barmherzige Brüder Regensburg Hospital.
+In this work, we present a preliminary study introducing a method to label key anatomical points, automatically process videos, and classify GMs based on the tracked anatomical points across videos, using a dataset from the Barmherzige Brüder Regensburg Hospital.
 The challenges posed by the variability in recording lengths and devices, the diversity of video scenarios (including both hospital and home environments), and the nature of the movement quality annotations—one per recording—are significant for reliable GM classification.
@@
 2 Dataset and Methods
-This work involves labelling specific anatomical landmarks of interest—so-called keypoints (KP)—in video frames (Section 2.2), tracking their positions across the video (Section 2.2), and extracting features from the tracked keypoints (Section 2.4) to classify movement quality.
-The full pipeline is illustrated in Figure 1.
+This work involves extracting features from specific anatomical landmarks of interest—so-called keypoints (KP)— in video frames (Section 2.4) to classify movement quality.
+We explore two means of acquiring the desired keypoints: 1) Label & Track, manually labelling specific KPs in a video frame (Section 2.2), and tracking their positions across the video (Section 2.2), as shown in the full pipeline illustrated in Figure 1, and 2) Automatic Label, automatically labeling the KPs in each frame, using the existing ML labeler AggPose Cao et al.
+(2022).
+The code [^1] https://github.com/chopardda/GMA is publicly available.
 Figure 1: Flowchart of keypoint extraction pipeline after initial video resampling and cropping.
@@
 There is also significant variability in camera angles and orientations, distance from subject to camera, infants’ clothing, and background types, among other factors.
-The data is split into two age-based groups: the early infancy group includes 39 infants with a postmenstrual age of up to 37 weeks, while the late infancy group consists of 37 infants with postmenstrual ages ranging from 46 to 56 weeks.
-Each sample is associated with a binary label that classifies the infant’s movement quality.
-In the early group, labels distinguish between normal and poor movement repertoire, while in the late group, labels indicate normal or absence of fidgety movements, following Einspieler and Prechtl (2005).
+The data is divided into two age-based groups according to the GM phase.
+The early General Movement group includes 39 infants with a postmenstrual age of 32 to 36 weeks (mean: 33 weeks), while the late General Movement group comprises of 37 infants with postmenstrual ages ranging from 49 to 59 weeks (mean: 53 weeks).
+Each sample is associated with a binary label reflecting the infant’s movement quality.
+Following Einspieler & Prechtl (2005), in the early GM group, which encompasses preterm and writhing GMs, labels distinguish between normal and poor movement repertoire.
+In contrast, the late GM group labels indicate the presence or absence of fidgety movements, with the presence of fidgety movements considered normal.
 To minimize bias and improve reliability, two trained physicians independently annotated the videos.
 In cases of disagreement, the recordings were reviewed jointly until a consensus label was reached.
-In the dataset, 24 infants in the early group exhibit poor repertoire, while 10 in the late group lack fidgety movements.
+In the dataset, 24 infants in the early GM group exhibit poor repertoire, while 10 in the late GM group lack fidgety movements.
+Further details are included in Appendix A.
 2.2 Keypoint Retrieval Tools
+2.2.1 Label & Track
 Keypoint Labelling
@@
 In this study, three independent non-expert labellers manually labelled all keypoints for subsequent tracking.
-Full lists of relevant keypoints can be found in Appendix B, see Figure 2 a) for an illustration of the tool.
+Full lists of relevant keypoints can be found in Figure 5, and  Figure 2 a) provides an illustration of the tool.
 Point Tracking
@@
 This approach predicts the trajectory of each point by extracting dense feature representations for every frame, leveraging a cost-volume approach.
-The cost volume measures the similarity between the feature of the tracked point in one frame and the features at all possible locations in subsequent frames, encoding the probability of the tracked point’s next position.
+Their model, TAP-Net, is designed to track arbitrary physical points on surfaces over video sequences.
+It is trained on a mix of synthetic and real-world datasets.
+The primary training datasets are synthetic, including TAP-Vid-Kubric, which contains simulated videos with physically realistic interactions between deformable and rigid objects, and TAP-Vid-RGB-Stacking, a robotic environment with teleoperated geometric shapes, focusing on challenges like tracking textureless and symmetric objects.
+These datasets provide perfect ground-truth point tracks, enabling TAP-Net to learn precise point tracking in complex, occlusion-prone scenarios.
+It uses a cost-volume-based architecture, inspired by optical flow methods, to track point motion across frames by comparing dense feature grids at query points.
+In particular, the cost volume measures the similarity between the feature of the tracked point in one frame and the features at all possible locations in subsequent frames, encoding the probability of the tracked point’s next position.
 A neural network then processes this cost volume to predict the tracked point’s location across frames.
-More details are provided in Appendix C.
-Figure 2: a) The point labelling tool prompts the user to label either only extreme keypoints or all keypoints.
+The network is trained with two objectives: (1) Point location regression, using Huber loss to estimate the query point’s position and ensure precise tracking, and (2) Occlusion classification, using cross-entropy loss to predict when the point is occluded, enabling the model to handle occlusion scenarios effectively.
+TAP-Net is validated and tested on real-world datasets like TAP-Vid-Kinetics, which includes diverse human actions and complex scenes, and TAP-Vid-DAVIS, which focuses on challenging segment and object tracking tasks.
+Figure 2: a) The manual point labelling tool prompts the user to label either only extreme keypoints or all keypoints.
 b) The outlier detection tool displays potential unrealistic point movements and allows manual correction.
+c) Compared to manual labelling, some keypoints are missing (head top, ears, and nose), while others are added (eyes, hand and foot extremities, torso)–displayed in grey–when labelling automatically with AggPose.
+2.2.2 AggPose: Automatic label
+In addition to manually labelled keypoints, we evaluate as a benchmark an automatic labelling approach using AggPose (Cao et al., 2022), a deep aggregation vision transformer designed specifically for infant pose estimation and trained on a large-scale infant pose dataset.
+Unlike traditional convolutional models, AggPose discards convolutional layers in favour of a fully transformer-based architecture, leveraging multi-scale feature aggregation through a deep-layer fusion mechanism.
+This method enhances spatial information sharing across different resolution levels, improving the robustness of KP detection in challenging infant pose scenarios.
+In this study, we use AggPose to automatically label KPs in each frame of the video dataset, extracting the full set of coordinates across time.
+This method encompasses a total of 21 KPs, slightly larger than the prior approach.
+See Figure 2 c) for a visualization of the KPs.
 2.3 Data Preprocessing
@@
 Outlier Detection
-To address inaccuracies in the tracking algorithm, such as when points jump inappropriately due to occlusions or overlapping limbs, an outlier detection mechanism is applied.
+To address inaccuracies in the tracking algorithm in the Label & Track (See 2.2.1) approach, such as when points jump inappropriately due to occlusions or overlapping limbs, an outlier detection mechanism is applied.
 The tool automatically flags any sudden and unrealistic movement of keypoints and prompts manual relabelling by the annotators, as shown in Figure 2 b).
+Specifically, we define such movements as coordinate changes that exceed 15 times the overall standard deviation for each keypoint.
+This threshold is set deliberately high to avoid flagging normal, natural movements, as well as fidgety movements that can inherently exhibit high magnitudes without indicating error.
 After relabelling, the affected keypoints are retracked.
-This process is iterated a fixed number of rounds, two in these results, or until no further outliers are detected, ensuring the accuracy of the tracked coordinates.
+This process is iterated a fixed number of rounds, two in these results, as we observed only marginal improvements in performance after three rounds in the final detection task, or until no further outliers are detected, ensuring the accuracy of the tracked coordinates.
+In general, this outlier-detection process can be structured to run for a set number of iterations, until a defined performance tolerance is reached, or until no further outliers are identified, ensuring the accuracy of the tracked coordinates.
 Final Data Preparation
@@
 2.4 Feature Extraction
-We consider two feature sets from the tracked keypoints: $x$- and $y$-coordinates of all keypoints and the angles between selected keypoints, based on Prakash et al.
-(2023) and further described in Appendix E.
+We consider three feature sets from the tracked keypoints to perform the classification task: $x$- and $y$-coordinates of all keypoints and the angles between selected keypoints, based on Prakash et al.
+(2023).
+Figure 3: List of all and extreme keypoints considered in this study for manual labelling, coloured by label on the right.
+All Keypoints Extreme
+nose
+head bottom
+head top x
+left ear
+right ear
+left shoulder
+right shoulder
+left elbow x
+right elbow x
+left wrist x
+right wrist x
+left hip
+right hip
+left knee x
+right knee x
+left ankle x
+right ankle x
+Figure 4: Example of frame with labelled keypoints.
+Figure 5: Illustration of the angle features. The dash lines show the angles that we compute in the set of angle features.
 Keypoint Coordinates
-The $x$- and $y$-coordinates of all labelled keypoints are treated independently as separate input channels, and used as a simple representation of the keypoint positions across frames.
+The $x$- and $y$-coordinates of all labelled keypoints are treated independently as separate input channels, and used as a simple representation of the keypoint positions across frames, as listed in Figure 5 and Figure 5.
 Keypoint Angles
 The angles are calculated between pairs of keypoints to capture movement dynamics.
-We compute the dot product of vectors, normalizing by magnitudes, and applying a clipping function for valid angle calculations in radians.
+We compute the dot product of vectors, normalizing by magnitudes, and applying a clipping function for valid angle calculations in radians, as illustrated in Figure 5.
+More precisely, given keypoints $\mathbf{p}_{1}$, $\mathbf{p}_{2}$, and $\mathbf{p}_{3}$ with coordinates $(x_{1},y_{1})$, $(x_{2},y_{2})$ and $(x_{3},y_{3})$ respectively, we would like to compute the angle between $\mathbf{p}_{1}$ and $\mathbf{p}_{3}$ at middle point $\mathbf{p}_{2}$.
+Thus, we define
+$\mathbf{v}_{1}=\begin{bmatrix}x_{1}-x_{2}\\ y_{1}-y_{2}\end{bmatrix},\quad\mathbf{v}_{2}=\begin{bmatrix}x_{3}-x_{2}\\ y_{3}-y_{2}\end{bmatrix}.$
+Then, the angle $\theta$ between the two vectors can be computed as:
+$\theta=\arccos\left(\frac{\mathbf{v}_{1}\cdot\mathbf{v}_{2}}{\|\mathbf{v}_{1}\|\|\mathbf{v}_{2}\|}\right),$
+where $\mathbf{v}_{1}\cdot\mathbf{v}_{2}$ is the dot product of the vectors and $\|\mathbf{v}_{1}\|$ and $\|\mathbf{v}_{2}\|$ are the magnitudes (or norms) of the vectors.
+Table 1 provides the list of the studied angle sets, which keypoints they are formed by, and a brief intuitive explanation.
+The angles are illustrated in 5.
+While there are more keypoints in the automatic labelling, the angle set is the same.
+$\mathbf{p}_{1}$ $\mathbf{p}_{2}$ $\mathbf{p}_{3}$ Explanation
+head top nose head bottom Head top to neck angle
+right ear nose left ear Right to left ear angle
+left elbow left shoulder head bottom Head to left shoulder angle
+right elbow right shoulder head bottom Head to right shoulder angle
+left wrist left elbow left shoulder Left elbow angle
+right wrist right elbow right shoulder Right elbow angle
+left knee left hip left shoulder Left hip angle
+right knee right hip right shoulder Right hip angle
+left hip left knee left ankle Left knee angle
+right hip right knee right ankle Right knee angle
+Table 1: List of angle components used in the study with intuitive explanations.
+Both The combination of both coordinates and angles, shown as ”both” in the results.
+2.5 Classification Models
+This work explores three different classification models for GMA, using the extracted time series described in Section 2.4.
+5-fold stratified cross-validation (CV) is performed across models to ensure robustness.
+1D Convolutional Neural Network (1D-CNN)
+1D-CNNs are a variant of the more common 2D convolutional neural network, which apply convolution operations to 1D signals, making them well-suited for processing time-series data.
+These networks have been shown to perform well when working with smaller labelled datasets and for specific applications (Kiranyaz et al., 2021).
+Long Short-Term Memory (LSTM)
+LSTM networks (Hochreiter & Schmidhuber, 1997) are a type of recurrent neural network capable of exploiting the temporal aspect of our data by learning relevant context information from previously seen time points.
+LSTMs are ideal for handling time series and assessing whether longer temporal memory helps in recognizing and classifying infant’s general movements.
+Random Forest (RF)
+RFs (Breiman, 2001) are powerful discriminative classifiers that perform well on a variety of datasets with minimal tuning (Biau & Scornet, 2016).
+It is an ensemble method that builds multiple decision trees trained on a random subset of data and combines their outputs to improve prediction accuracy and reduce overfitting.
+Although they do not explicitly account for the temporal structure of the data, their efficiency and accuracy make them a strong candidate for our classification tasks.
+3 Experiments and Results
+3.1 Implementation details
+We performed a grid search over the hyperparameter space for each of the three classifier types in Section 2.5, choosing the values that performed best in the evaluation.
+The final models were trained using these hyperparameters with 5-fold stratified CV, and the results from the held-out test set, with 20% of the data, were collected.
+On the Label & Track approach, the classification models used included a 1D-CNN with binary cross entropy (BCE) loss, learning rate of 0.00001, batch size 6, a fully-connected bottleneck with 150 features, and trained for 150 epochs.
+The LSTM used the BCE loss, learning rate 0.001, batch size 6, 3 layers, a hidden size of 64 and trained for 200 epochs.
+Finally, the random forest had 170 estimators.
+Moreover, using the Aggpose approach, the 1D-CNN had a learning rate of 0.0001, batch size of 4 and 150 features in the fully connected bottleck trained for 50 epochs.
+The LSTM also used BCE loss learning rate 0.0001, batch size 4, 2 layers, a hidden size of 64 and trained for 50 epochs.
+Lastly, the random forest in this setup had 45 estimators.
+All experiments were run using 30 independent seeds on each individual set.
+Our models were implemented using pytorch v2.3.1 and scikit-learn v1.3.0.
+The final classification was evaluated using accuracy, Area Under the Receiver Operating Characteristic Curve (AUROC), and Area Under the Precision–Recall curve (AUPRC).
+3.2 Results
 Table 2: AUROC of various classifiers for GM classification across different keypoint labellers.
@@
 Best result per model and labeller is bolded.
-Keypoint Labeller Input Features Early Infancy Group AUROC Late Infancy Group AUROC
+Keypoint Labeller Input Features Early GM Group AUROC Late GM Group AUROC
 RF LSTM CNN RF LSTM CNN
@@
 Both $0.6964$ $0.6454$ 0.7484 $0.4771$ $0.4747$ $0.5550$
-2.5 Classification Models
-This work explores three different classification models for GMA, using the extracted time series described in Section 2.4.
-5-fold stratified cross-validation (CV) is performed across models to ensure robustness.
-1D Convolutional Neural Network (1D-CNN)
-1D-CNNs are a variant of the more common 2D convolutional neural network, which apply convolution operations to 1D signals, making them well-suited for processing time-series data.
-These networks have been shown to perform well when working with smaller labelled datasets and for specific applications (Kiranyaz et al., 2021).
-Long Short-Term Memory (LSTM)
-LSTM networks (Hochreiter and Schmidhuber, 1997) are a type of recurrent neural network capable of exploiting the temporal aspect of our data by learning relevant context information from previously seen time points.
-LSTMs are ideal for handling time series and assessing whether longer temporal memory helps in recognizing and classifying infant’s general movements.
-Random Forest (RF)
-RFs (Breiman, 2001) are powerful discriminative classifiers that perform well on a variety of datasets with minimal tuning (Biau and Scornet, 2016).
-It is an ensemble method that builds multiple decision trees trained on a random subset of data and combines their outputs to improve prediction accuracy and reduce overfitting.
-Although they do not explicitly account for the temporal structure of the data, their efficiency and accuracy make them a strong candidate for our classification tasks.
-3 Experiments and Results
-3.1 Implementation details
-We performed a grid search over the hyperparameter space for each of the three classifier types in Section 2.5, choosing the values that performed best in the evaluation.
-The final models were trained using these hyperparameters with 5-fold stratified CV, and the results from the held-out test set, with 20% of the data, were collected.
-The classification models used included a 1D-CNN with binary cross entropy (BCE) loss, learning rate of 0.00001, batch size 6, a fully-connected bottleneck with 150 features, and trained for 150 epochs.
-The LSTM used the BCE loss, learning rate 0.001, batch size 6, 3 layers, a hidden size of 64 and trained for 200 epochs.
-Finally, the random forest had 170 estimators.
-All experiments were run using 30 independent seeds on each individual set.
-Our models were implemented using pytorch v2.3.1 and scikit-learn v1.3.0.
-The final classification was evaluated using accuracy, Area Under the Receiver Operating Characteristic Curve (AUROC), and Area Under the Precision–Recall curve (AUPRC).
-3.2 Results
-Table 1 shows the results of the different classification models, for each labeller individually as well as averaged, measured in AUROC.
-See Appendix F for a comprehensive overview of AUPRC and accuracy in each task.
+Coord. $0.4780\pm 0.19$ $0.5721\pm 0.19$ $0.4643\pm 0.18$ $0.4001\pm 0.26$ $0.5436\pm 0.23$ $0.5628\pm 0.24$
+Automatic Angles $\textbf{0.5076}\pm 0.19$ $\textbf{0.6112}\pm 0.21$ $0.4903\pm 0.15$ $\textbf{0.4381}\pm 0.25$ $\textbf{0.6586}\pm 0.24$ $\textbf{0.7156}\pm 0.24$
+Both $0.4819\pm 0.17$ $0.5330\pm 0.21$ $\textbf{0.5101}\pm 0.19$ $0.4036\pm 0.24$ $0.6483\pm 0.26$ $0.5208\pm 0.29$
+Table 2 shows the results of the different classification models, for each labeller individually as well as averaged and labeled with AggPose, measured in AUROC.
+See Appendix B for a comprehensive overview of AUPRC and accuracy in each task.
 They demonstrate a variety of outcomes across age groups and feature sets, with no clear pattern indicating that one method is superior.
 However, the use of angle features, either isolated or combined, proves useful.
-Additionally, the early group shows significantly better overall results, showing a big difference in the difficulty of the task across age groups.
+Additionally, the early GM group shows significantly better overall results, showing a big difference in the difficulty of the task across age groups.
+Overall, the Label & Track approach performs comparably to the automatically labeled AggPose method.
+However, in a significant number of data groups, it demonstrates improved results, highlighting the positive impact of keypoint tracking.
+By incorporating temporal information from the time series, this approach enhances classification performance, whereas AggPose labels each keypoint in isolated frames, missing potential contextual cues.
 One of the key takeaways from this study is that despite the relatively small and heterogeneous dataset, our approach has shown success in automatically classifying GM in newborns, with average AUROCs of up to 0.8283 in certain scenarios, opening promising lines of research for GMA “in the wild”.
 4 Conclusion and Future Outlook
-In conclusion, our study demonstrates that we can successfully predict GMA with a promising AUROC of up to 0.8283, with performance generally better for the early group.
+In conclusion, our study demonstrates that we can successfully predict GMA with a promising AUROC of up to 0.8283, with performance generally better for the early GM group.
 This highlights the potential of using automated methods to assess infant motor behaviour.
@@
 This study lays the groundwork for further advancements, emphasizing the need for larger datasets and more refined models and feature extraction approaches to capture the data complexities.
-Ultimately, the goal is to enhance early prediction of neurodevelopmental diseases, offering a valuable tool for clinicians and ultimately improving patient care.
+Ultimately, the goal is to enhance early prediction of neurodevelopmental diseases, offering a valuable tool for clinicians, and ultimately improving patient care.
 One of the key weaknesses of this study is the relatively small dataset, which limits the robustness and generalizability of our results, as indicated by relatively large standard deviations across folds.
@@
 Expanding the dataset would allow us to explore more sophisticated methods, such as Multiple Instance Learning (MIL), which could introduce interpretability by making predictions on a per-episode basis.
-Additionally, using Vision Transformers to directly process the videos could better capture spatial and temporal patterns, further improving classification accuracy.
+Similarly, using Vision Transformers to directly process the videos could help better capture spatial and temporal patterns, further improving classification performance.
 Future work within the current dataset could focus on developing further features from the extracted keypoints and exploring other model classes to leverage the temporal structure in the task.
-[…]
+Acknowledgments
 DC is funded from the Strategic Focal Area “Personalized Health and Related Technologies (PHRT)” grant #2021-911 of the ETH Domain (Swiss Federal Institutes of Technology), and SL and KC from the Swiss State Secretariat for Education, Research and Innovation (SERI), contract #MB22.00047.
@@
 Appendix A Extended Dataset Description
-In this section we present additional details about our dataset to the best of our ability.
-Some details are not available to us, such as weight at video for early group infants.
-Table 2 shows several demographic characteristics, as well as details about the videos themselves for both the early and late groups.
-Figure 3 provides histograms for the distributions of both video resolution and FPS for the early and late video groups.
+In this section, we present additional details about our dataset to the best of our ability.
+Some details are not available to us, such as weight at video reconding time for early GM group infants.
+Table 3 shows several demographic characteristics, as well as details about the videos themselves for both the early and late GM groups.
+Figure 6 provides histograms for the distributions of both video resolution and FPS for the early and late GM groups.
 Table 3: Calculated demographic and video characteristics
-Characteristic Early Group Late Group
-Age at video recording (Days) (mean, stddev) 30.9, 19.9 170.4, 24.8
-Post-menstrual age at video recording (Days) (mean, stddev) 231.5, 6.5 N/A
-Corrected Age at video recording (Days) (mean, stddev) N/A 92.5, 17.4
-Weight at video recording (g) (mean, stddev) N/A 4148.2, 830.7
-Video Length (s) (mean, stddev) 409, 144 224, 93
-FPS (min, max) 30, 120 30, 120
+Characteristic Early GM Group Late GM Group
+Age at video recording (Days) (mean $\pm$ stddev) $30.9\pm 19.9$ 170.4 $\pm$ 24.8
+Post-menstrual age at video recording (Days) (mean $\pm$ stddev) 231.5 $\pm$ 6.5 N/A
+Corrected Age at video recording (Days) (mean $\pm$ stddev) N/A 92.5 $\pm$ 17.4
+Weight at video recording [g] (mean, stddev) N/A 4148.2 $\pm$ 830.7
+Video Length [s] (mean $\pm$ stddev) 409 $\pm$ 144 224 $\pm$ 93
+FPS [min, max] [30, 120] [30, 120]
 Extremely Preterm 16 15
@@
 Not Preterm 0 0
-Figure 3: Histograms showing the distributions of a) early infancy group video resolution, b) early infancy group video FPS, c) late infancy group video resolution, and d) late infancy group video FPS
-Appendix B Keypoint Labelling Tool
-Figure 5 illustrates the keypoints used in this study and the labelling tool developed.
-Figure 4:
-List of all and extreme keypoints considered in this study, coloured by label on the right.
-All Keypoints Extreme nose head bottom head top x left ear right ear left shoulder right shoulder left elbow x right elbow x left wrist x right wrist x left hip right hip left knee x right knee x left ankle x right ankle x
-Figure 5: Example of frame with labelled keypoints.
-Appendix C Keypoint Tracking
-We adopt the work by Doersch et al.
-(2022) for keypoint tracking.
-Their model, TAP-Net, is designed to track arbitrary physical points on surfaces over video sequences.
-It is trained on a mix of synthetic and real-world datasets.
-The primary training datasets are synthetic, including TAP-Vid-Kubric, which contains simulated videos with physically realistic interactions between deformable and rigid objects, and TAP-Vid-RGB-Stacking, a robotic environment with teleoperated geometric shapes, focusing on challenges like tracking textureless and symmetric objects.
-These datasets provide perfect ground-truth point tracks, enabling TAP-Net to learn precise point tracking in complex, occlusion-prone scenarios.
-It uses a cost-volume-based architecture, inspired by optical flow methods, to track point motion across frames by comparing dense feature grids at query points.
-The network is trained with two objectives: (1) Point location regression, using Huber loss to estimate the query point’s position and ensure precise tracking, and (2) Occlusion classification, using cross-entropy loss to predict when the point is occluded, enabling the model to handle occlusion scenarios effectively.
-TAP-Net is validated and tested on real-world datasets like TAP-Vid-Kinetics, which includes diverse human actions and complex scenes, and TAP-Vid-DAVIS, which focuses on challenging segment and object tracking tasks.
-Appendix D Outlier Detection
-To mitigate potential errors from the tracking tool, we implemented a tool that automatically flags any sudden, unrealistic movements of keypoints.
-Specifically, we define such movements as coordinate changes that exceed 15 times the overall standard deviation for each keypoint.
-This threshold is set deliberately high to avoid flagging normal, natural movements, as well as fidgety movements that can inherently exhibit high magnitudes without indicating error.
-Once flagged, the affected keypoints are relabeled and subsequently retracked.
-This process is iteratively applied up to three times, as we observed only marginal improvements in performance after three rounds in the final detection task.
-In general, this outlier-detection process can be structured to run for a set number of iterations, until a defined performance tolerance is reached, or until no further outliers are identified, ensuring the accuracy of the tracked coordinates.
-Appendix E Feature Extraction
-In this work, we perform the classification task on a set of features extracted from the obtained keypoints. In particular, we study the classification on three subsets, as reported in our results:
-  • Coord. The keypoint coordinates, that correspond to the x- and y-locations of all keypoints, as listed in Figure 5.
-  • Angles The angles between keypoints of interest. The values of the angle corresponding to each keypoint pair are provided separately to the model.
-    $\mathbf{p}_{1}$ $\mathbf{p}_{2}$ $\mathbf{p}_{3}$ Explanation
-    head top nose head bottom Head top to neck angle
-    right ear nose left ear Right to left ear angle
-    left elbow left shoulder head bottom Head to left shoulder angle
-    right elbow right shoulder head bottom Head to right shoulder angle
-    left wrist left elbow left shoulder Left elbow angle
-    right wrist right elbow right shoulder Right elbow angle
-    left knee left hip left shoulder Left hip angle
-    right knee right hip right shoulder Right hip angle
-    left hip left knee left ankle Left knee angle
-    right hip right knee right ankle Right knee angle
-    Table 3: List of angle components used in the study with intuitive explanations.
-    Figure 6: Illustration of the angle features. The dash lines show the angles that we compute in the set of angle features.
-    Given keypoints $\mathbf{p}_{1}$, $\mathbf{p}_{2}$, and $\mathbf{p}_{3}$ with coordinates $(x_{1},y_{1})$, $(x_{2},y_{2})$ and $(x_{3},y_{3})$ respectively.
-We would like to compute the angle between $\mathbf{p}_{1}$ and $\mathbf{p}_{3}$ at middle point $\mathbf{p}_{2}$.
-Thus, we define
-    $\mathbf{v}_{1}=\begin{bmatrix}x_{1}-x_{2}\\ y_{1}-y_{2}\end{bmatrix},\quad\mathbf{v}_{2}=\begin{bmatrix}x_{3}-x_{2}\\ y_{3}-y_{2}\end{bmatrix}.$
-    Then, the angle $\theta$ between the two vectors can be computed as:
-    $\theta=\arccos\left(\frac{\mathbf{v}_{1}\cdot\mathbf{v}_{2}}{\|\mathbf{v}_{1}\|\|\mathbf{v}_{2}\|}\right),$
-    where $\mathbf{v}_{1}\cdot\mathbf{v}_{2}$ is the dot product of the vectors and $\|\mathbf{v}_{1}\|$ and $\|\mathbf{v}_{2}\|$ are the magnitudes (or norms) of the vectors.
-    Table 3 provides the list of the studied angle sets, which keypoints they are formed by, and a brief intuitive explanation. The angles are illustrated in 6.
-  • Both The combination of both coordinates and angles, shown as ”both” in the results.
-Appendix F Further results
+Figure 6: Histograms showing the distributions of a) early GM group video resolution, b) early GM group video FPS, c) late GM group video resolution, and d) late GM group video FPS
+Appendix B Further results
 Table 4 and Table 5 show the results of the presented classification models, for each labeller individually as well as averaged, grouped by age and feature extraction method used, measured in AUPRC and accuracy, respectively.
@@
 Table 4: Results. AUPRC over 30 seeds is reported. The best result for each model and labeller is bolded.
-Labeller Features Early Infancy Group Late Infancy Group
+Labeller Features Early GM Group Late GM Group
 AUPRC AUPRC
@@
 Both $0.7928$ $0.7308$ $0.8360$ $0.3335$ $0.3527$ $0.4297$
+Coord. $0.6324\pm 0.16$ $0.6969\pm 0.15$ $0.65\pm 0.15$ $0.3560\pm 0.21$ $0.5015\pm 0.24$ $0.4973\pm 0.25$
+Automatic Angles $\textbf{0.6792}\pm 0.15$ $\textbf{0.7323}\pm 0.16$ $0.6546\pm 0.13$ $\textbf{0.3620}\pm 0.19$ $0.5765\pm 0.27$ $\textbf{0.6523}\pm 0.28$
+Both $0.6340\pm 0.16$ $0.6451\pm 0.16$ $\textbf{0.6769}\pm 0.15$ $0.3340\pm 0.18$ $\textbf{0.5923}\pm 0.28$ $0.4504\pm 0.28$
 Table 5: Results. Accuracy over 30 seeds is reported. The best result for each model and labeller is bolded.
-Labeller Features Early Infancy Group Late Infancy Group
+Labeller Features Early GM Group Late GM Group
 Accuracy Accuracy
@@
 Both $0.6017$ 0.6272 0.6742 $0.6374$ $0.6244$ 0.6225
+Coord. $0.4679\pm 0.13$ $0.5382\pm 0.16$ $0.4663\pm 0.15$ $0.5393\pm 0.17$ $0.599\pm 0.17$ $0.6236\pm 0.18$
+Automatic Angles $\textbf{0.5015}\pm 0.17$ $\textbf{0.5907}\pm 0.14$ $0.4707\pm 0.11$ $\textbf{0.6185}\pm 0.11$ $\textbf{0.7131}\pm 0.15$ $\textbf{0.7048}\pm 0.18$
+Both $0.4745\pm 0.15$ $0.5285\pm 0.20$ $\textbf{0.4879}\pm 0.14$ $0.5948\pm 0.14$ $0.6101\pm 0.18$ $0.6127\pm 0.19$